# AfroXLMR-Social: Adapting Pre-trained Language Models for African Languages Social Media Text

### v2

@@ 6.2 Task Adaptive Pretraining (TAPT)
Table 4: Cross-TAPT results across the three datasets. Base column is baseline results from AfroXLMR, DATPT + TAPT column results for AfriSenti and AfriHate are TAPT from the AfriEmo dataset. DATPT + TAPT for AfriEmo results is TAPT from the AfriSenti dataset. Reported results are the Macro F1 score. Blank values (-) indicate that the specific dataset does not cover the language.
@@
Results Table 4 shows the TAPT results across datasets and languages. As a result, all TAPT performs better than the base model and equal or competitive results with the DAPT. This indicates we can achieve better results for our targeted evaluation task using very small, quality, and task-related data. In our case, using AfriSenti data without the labels as a TAPT is very helpful for the fine-grained multi-label emotion classification task. For the AfriSenti sentiment evaluation task, TAPT with AfriHate data has better average results than TAPT with AfriEmo. For the AfriHate evaluation, TAPT in AfriSenti data has better average results than TAPT with AfriEmo. In addition to the task similarity, this improvement might be affected by the number of total instances in each dataset (the total instances of AfriSenti is 107,694, AfriEmo 70,859, and AfriHate 90,455) and the vocabulary similarity across the datasets.
@@ 6.3 Combining DAPT and TAPT
We explore the effect of both adaptation techniques by combining DAPT and TAPT. In this approach, we apply DAPT to the base model and then the TAPT training. These phases of pretraining add up to make this approach the more computationally expensive setting.
Results The results are shown in Table 4, DAPT + TAPT column. The results show that the subjective NLP tasks benefit from the combined DAPT and TAPT approaches. DAPT followed by TAPT achieves the best performance. However, first adapting the model to the domain (DAPT), then applying TAPT would be susceptible to catastrophic forgetting of the domain-relevant corpus Yogatama et al. (2019); alternate methods of combining the procedures may result in better downstream performance. This is shown from the summarized result Table 5 that sometimes the DATP + TAPT result performs less than the DAPT-only results, while it is better than the baseline results.
@@
6.4 AfroXLMR-Social Vs. LLMs
Figure 3: AfriEmo Macro F1 results from AfroXLMR-Social and LLMs. The results for LLMs are based on zero-shot evaluations, selecting the best results from five different types of prompts. The benchmark results for the LLMs are taken from the SemEval-2025 Task 11 dataset papers Muhammad et al. (2025b); Belay et al. (2025b).
@@ 7 Conclusion
This work explores the effects of the domain adaptive pretraining (DAPT) and task adaptive pertaining (TAPT) across three subjective tasks of African languages. We created the AfriSocial corpus, a social domain-specific corpus sourced from X and news. Using AfriSocial, we further developed the AfroXLMR-Social language model, which specialized in the social domain. We improve the performance of evaluated tasks, sentiment analysis (AfriSenti), emotion analysis (AfriEmo), and hate speech classification (AfriHate) using DAPT and TAPT techniques. We show that pretraining the model towards a small domain-specific corpus and related task-relevant data can provide significant improvements. While the combination of the two methods, DAPT + TAPT, also achieves better results than baseline models, TAPT followed by DAPT would be susceptible to catastrophic forgetting of the task-relevant corpus. We achieve better state-of-the-art results using a small domain-related corpus from the encoder-only model than state-of-the-art large-language models (LLMs). AfriSocial and AfroXLMR-Social will support the development of African languages in the NLP and improve similar-sourced tasks. It opens further domain explorations as the AfriSocial X and news domains are also available separately.
@@
This work explores the social media X data and news on downstream more subjective tasks such as sentiment analysis, emotion analysis, and hate speech classification. Other outside-the-domain data, such as the health domain corpus on social media-based tasks and other cross-domain evaluations, are another interesting way to explore. We have used a total of 3.5M sentences for the domain adaptation (DAPT) exploration, with a considerable variation in data statistics among languages. More balanced data from the sources across languages might provide more insights into the evaluation.
@@
In this work, our evaluation is limited to three subjective tasks: sentiment analysis, emotion analysis, and hate speech classification, as we aimed to show the effects of DATP and TAPT approaches on the social domain. This work can be extended using the proposed approaches and the AfriSocial domain-specific pre-training corpus to evaluate other similar tasks.
@@ Appendix D Training Details of DAFT and TAPT
Hyper-parameters for adapters We trained the task adapter using the following hyper-parameters: batch size of 32, 10 epochs, and learning rate of 5e-5. For TAPT, the parameters are similar to those of DAPT, except that the batch size is 8. We used their tokenizer, for XLMR - XLMR tokenizer, and AfroXLMR - AfroXLMR tokenizer. PyTorch was used for fine-tuning, and pre-trained models were sourced from Hugging Face. The domain adaptive fine-tuning training is trained on three distributed-based GPUs. Following standard practice, we pass the final layer [CLS] token representation to a task-specific feedforward layer for prediction with three epochs. The reported results from fine-tuned pre-trained language models are the average results of five runs.
@@ E.1 Encoder-only Langauge Models
  • XLM-RoBERTa Conneau et al. (2020b) - FacebookAI/xlm-roberta-base (large)
@@
  • AfroXLMR Alabi et al. (2022) - Davlan/afro-xlmr-large-61L (76L)
@@ E.2 Open-source LLMs
  • Llama 3 8B Grattafiori et al. (2024) - meta-llama/Meta-Llama-3-8B-Instruct
  • Llama 3.1 (8B, 70B) Grattafiori et al. (2024) - meta-llama/Llama-3.1-8B-Instruct and meta-llama/Llama-3.1-70B-Instruct, respectively.
@@
  • Gemma 2 (9B, 27B) Team et al. (2024) - google/gemma-2-2b-it and google/gemma-2-27b-it
  • DeepSeek-R1-70 Guo et al. (2025) - deepseek-ai/DeepSeek-R1-Distill-Llama-70B
@@ E.3 Propritary LLMs
  • Gemini 1.5 Pro Georgiev et al. (2024) - Gemini 1.5 Pro 002 accessed via Google API

### v3

@@ 4.3 AfriHate: Hate Speech Dataset
Table 4: Cross-TAPT results across the three datasets. Base column is baseline results from AfroXLMR, DATPT + TAPT column results for AfriSenti and AfriHate are TAPT from the AfriEmo dataset. DATPT + TAPT for AfriEmo results is TAPT from the AfriSenti dataset. Reported results are the Macro F1 score. Blank values (—) indicate that the specific dataset does not cover the language.
5 Language Models
5.1 Encoder-only Language Models
Multilingual encoder-only pretrained language models (PLMs) such as XLM-R Conneau et al. (2020) and mBERT Devlin et al. (2019) have shown impressive capability on many languages for a variety of downstream tasks. They are also often used to initialize checkpoints to adapt to other languages, such as AfroXLMR-76L Adelani et al. (2024), which is initialized from XLM-R to specialize in African languages. We evaluate popular multilingual and African-centric PLMs such as AfroLM Dossou et al. (2022) and AfriBERTa Ogueji et al. (2021) and found that AfroXLMR is better for the targeted evaluation datasets as it covers more African languages. We make AfroXLMR our benchmarks and further train on the AfriSocial domain and the selected evaluation tasks. The AfroXLMR followed by DAPT with the AfriSocial domain-specific corpus gives us AfroXLMR-Social. The detailed hyperparameters of the continual training are shown in Appendix D.
5.2 Large Language Models
We compare our DAPT and TAPT approach results from AfroXLMR with state-of-the-art open source LLMs such as Llama 3 Dubey et al. (2024), Gemma 2 Riviere et al. (2024), Mistral Jiang et al. (2024), and proprietary LLMs such as Gemini 1.5 Reid et al. (2024) and GPT-4o Hurst et al. (2024). For AfriSenti and AfriHate task results, we use the LLMs benchmark results from the AfroBench Ojo et al. (2025) leaderboard. For the AfriEmo task, we use LLM results from the SemEval-2025 Task 11 datasets papers Muhammad et al. (2025b); Belay et al. (2025b). The detailed versions of the LLMs are presented in the Appendix E.
Figure 3: AfriEmo Macro F1 results from AfroXLMR-Social and LLMs. The results for LLMs are based on zero-shot evaluations, selecting the best results from five different types of prompts. The benchmark results for the LLMs are taken from the SemEval-2025 Task 11 dataset papers Muhammad et al. (2025b); Belay et al. (2025c).
6 Experiment Results
6.1 Domain Adaptive Pre-training (DAPT)
The domain-adaptive pre-training (DAPT) approach is straightforward; we continue pre-training from the strong baseline language model (AfroXLMR) using the domain-specific AfriSocial corpus in a multilingual setup. Baseline results from (AfroXLMR) and after applying DAPT are presented in Table 3.
Baseline: As a baseline, we evaluate various encoder-only models and found that AfroXLMR, which covers 76 African languages Alabi et al. (2022), performs better than other BERT-based encoder-only models across targeted datasets since it includes more African languages during pre-training. The evaluation results of othe encoder-only multilingual and African-centric models are shown in Appendix C.
Results: The results before and after DAPT are shown in Table 3, AfroXLMR and DAPT columns, respectively. We observe that DAPT improves over the baseline in almost all languages and datasets, demonstrating the benefit of DAPT when the target domain is relevant. It consistently improves over the baseline models for each language. It suggests that continual pre-training on a small, quality, domain-relevant dataset is important for subjective tasks from the same domain.
6.2 Task Adaptive Pre-training (TAPT)
Similar to DAPT, TAPT consists of a second phase of continual pre-training. TAPT is a cross-task transfer across datasets, which refers to finetuning on the unlabeled data of the non-targeted task during evaluation. We explore the TAPT approach by directly applying it to the base model followed by DAPT. For example, if we make TAPT for the AfriSenti task, we further fine-tune the base model and DAPT model using the unlabeled data of the AfriEmo and AfriHate datasets separately. Compared to DAPT, the task-adaptive approach strikes a different trade-off: it uses a far smaller pre-training corpus, assuming the training set represents aspects of the target task. This makes TAPT much less expensive to run than DAPT. When we apply TAPT across tasks, for example, TAPT with AfriSenti for AfriHate evaluation, we ensure that we exclude any duplication from the training data (in this case, AfriSenti) to prevent data leakage.
@@ 6.2 Task Adaptive Pre-training (TAPT)
Results Table 4 shows the TAPT results across datasets and languages. As a result, all TAPT performs better than the base model and equal or competitive results with the DAPT. This indicates we can achieve better results for our targeted evaluation task using very small, quality, and task-related data. In our case, using AfriSenti data without the labels as a TAPT is very helpful for the fine-grained multi-label emotion classification task. For the AfriSenti sentiment evaluation task, TAPT with AfriHate data achieves a better average results than TAPT with AfriEmo. For the AfriHate evaluation, TAPT in AfriSenti data has better average results than TAPT with AfriEmo. In addition to the task similarity, this improvement might be affected by the number of total instances in each dataset (the total instances of AfriSenti is 107,694, AfriEmo 70,859, and AfriHate 90,455) and the vocabulary similarity across the datasets.
@@ 6.3 Combining DAPT and TAPT
We explore the effect of both adaptation techniques by combining DAPT and TAPT. In this approach, we apply DAPT to the base model and then the TAPT training. These phases of pre-training add up to make this approach the more computationally expensive setting.
Results The results are shown in Table 4, DAPT + TAPT column. The results show that the subjective NLP tasks benefit from the combined DAPT and TAPT approaches. DAPT followed by TAPT achieves the best performance. However, first adapting the model to the domain (DAPT), then applying TAPT would be susceptible to catastrophic forgetting of the domain-relevant corpus Yogatama et al. (2019); alternate methods of combining the procedures may result in better downstream performance. This is shown from the summarized result in Table 5 that sometimes the DAPT + TAPT result performs less than the DAPT-only results, while it is better than the baseline results.
6.4 AfroXLMR-Social Vs. LLMs
@@ 7 Conclusion
This work explored the effects of domain-adaptive pre-training (DAPT) and task-adaptive pre-training (TAPT) across three subjective tasks involving African languages. We created the AfriSocial corpus, a social domain-specific corpus sourced from X and news. Using AfriSocial, we further developed the AfroXLMR-Social language model, which specialized in the social domain. We improved the performance of evaluated tasks, sentiment analysis (AfriSenti), emotion analysis (AfriEmo), and hate speech classification (AfriHate) using DAPT and TAPT techniques. We showed that pre-training the model towards a small domain-specific corpus and related task-relevant data can provide significant improvements. While the combination of the two methods, DAPT + TAPT, also achieved better results than the baseline models, TAPT followed by DAPT would be susceptible to catastrophic forgetting of the task-relevant corpus. We achieved better state-of-the-art results using a small domain-related corpus from the encoder-only model than state-of-the-art large-language models (LLMs). AfriSocial and AfroXLMR-Social will support the development of African languages in the NLP and improve similar-sourced tasks. It opened further domain explorations as the AfriSocial X and news domains are also available separately.
@@
This work focuses on social media data from X and news sources for downstream tasks that are inherently subjective, such as sentiment analysis, emotion recognition, and hate speech classification. Extending the evaluation to out-of-domain data (e.g., health forums, long-form blogs) and the impacts of topic variations (e.g., politics, sports, business, health) presents another promising avenue for understanding cross-domain generalization in social media–based tasks. For the domain-adaptive pre-training (DAPT) exploration, we utilized a corpus of 3.5M sentences, which exhibits substantial variation in data statistics across different languages.
@@
In this work, we restrict our evaluation to three subjective tasks—sentiment analysis, emotion recognition, and hate speech classification—in order to highlight the effects of DAPT and TAPT approaches within the social domain. Future work could extend these approaches to a broader range of downstream NLP tasks, including more knowledge-intensive and objective benchmarks such as question answering and machine translation, thereby offering a more comprehensive understanding of their generalizability and impact.
Evaluation of LLMs
Assessing the impact of DAPT and TAPT approaches on the latest LLM families—such as Llama, Gemini, GPT, Mistral, and others remains an open direction for future research. In-context learning evaluations, particularly in few-shot settings, provide a promising lens for understanding model behavior, while prompting strategies such as Chain-of-Thought (CoT) reasoning and in-domain prompting have demonstrated notable improvements in LLM performance across various tasks. Systematic evaluation of these techniques in combination with DAPT and TAPT may therefore yield more profound insights and potentially lead to different conclusions regarding the effectiveness and generalizability of such adaptation methods.
Imbalanced Data Across Languages.
As illustrated in the AfriSocial corpus (Table 2), there exists substantial variability in the availability of domain-specific data across languages (e.g., 8.6K sentences for Twi versus 866K for Zulu). Investigating the impact of such imbalances on the effectiveness of DAPT and TAPT continual pre-training approaches could yield valuable insights into both the robustness of adaptation techniques and language-specific behaviors. Incorporating more balanced data across languages in future work may further enhance the evaluation and provide a clear understanding of the dynamics of cross-lingual adaptation.
Acknowledgments
The work was done with partial support from the Mexican Government through the grant A1-S-47854 of CONACYT, Mexico, grants 20241816, 20241819, and 20240951 of the Secretaría de Investigación y Posgrado of the Instituto Politécnico Nacional, Mexico. The authors thank the CONACYT for the computing resources brought to them through the Plataforma de Aprendizaje Profundo para Tecnologías del Lenguaje of the Laboratorio de Supercómputo of the INAOE, Mexico and acknowledge the support of Microsoft through the Microsoft Latin America PhD Award.
Shamsuddeen acknowledges support from a Research Grant under the Nigeria Artificial Intelligence Research (NAIR) Scheme, administered by the National Information Technology Development Agency (NITDA), for developing a dataset for Nigerian languages.
@@ Appendix D Training Details of DAFT and TAPT
Hyper-parameters for adapters We trained the task adapter using the following hyper-parameters: batch size of 32, 10 epochs, and learning rate of 5e-5. For TAPT, the parameters are similar to those of DAPT, except that the batch size is 8. We used their tokenizer, for XLMR - XLMR tokenizer, and AfroXLMR - AfroXLMR tokenizer. PyTorch was used for fine-tuning, and pre-trained models were sourced from Hugging Face. The domain-adaptive fine-tuning training is trained on three distributed GPUs for 3 days, whereas TAPT finishes in less than one hour. Following standard practice, we pass the final layer [CLS] token representation to a task-specific feedforward layer for prediction with three epochs. The reported results from fine-tuned pre-trained language models are the average results of five runs.
@@ E.1 Encoder-only Langauge Models
  • XLM-RoBERTa Conneau et al. (2020) - FacebookAI/xlm-roberta-base (large)
@@
  • AfroXLMR Adelani et al. (2024) - Davlan/afro-xlmr-large-61L (76L)
@@ E.2 Open-source LLMs
  • Llama 3 8B Dubey et al. (2024) - meta-llama/Meta-Llama-3-8B-Instruct
  • Llama 3.1 (8B, 70B) Dubey et al. (2024) - meta-llama/Llama-3.1-8B-Instruct and meta-llama/Llama-3.1-70B-Instruct, respectively.
@@
  • Gemma 2 (9B, 27B) Riviere et al. (2024) - google/gemma-2-2b-it and google/gemma-2-27b-it
  • DeepSeek-R1-70 DeepSeek-AI et al. (2025) - deepseek-ai/DeepSeek-R1-Distill-Llama-70B
@@ E.3 Propritary LLMs
  • Gemini 1.5 Pro Reid et al. (2024) - Gemini 1.5 Pro 002 accessed via Google API